# A Robust Mixed-Effects Bandit Algorithm for Assessing Mobile Health Interventions

### v3

@@ 6 Experiments
This section presents results from applying DML-TS-NNR in a simulated mHealth study with staggered recruitment and an off-policy comparison study for cardiac rehabilitation mHealth interventions. The simulations were implemented using Python and the results were generated using individual compute nodes with two 3.0 GHz Intel Xeon Gold 6154 processors and 180 GB of RAM. Case studies were implemented using R 4.2.2 and results were generated on a cluster composed of individual compute nodes with 2.10 GHz Intel Xeon Gold 6230 processors and 192 GB of RAM.
@@ 6.1 Competitor Comparison Simulation
In this section, we compare DML-TS-NNR to four competing methods in simulation. We implemented DML-TS-NNR using Option 2 with a bagged ensemble of stochastic gradient trees (Gouk et al., 2019; Mastelini et al., 2021) trained online via the River library (Montiel et al., 2021). We programmed the linear algebra operations using SuiteSparse (Davis & Hu, 2011) to take advantage of the sparse nature of the Laplacian matrices and $\phi_{i,t}$ from Algorithm 2.
@@
We compare these methods under three settings: Homogeneous Users, Heterogeneous Users, and Nonlinear. The first two involve a linear baseline model and time-homogeneous parameters, with the second setting having distinct user parameters. The third setting, designed to mirror the challenges of mHealth studies, includes a nonlinear baseline and both user- and time-specific parameters. For each setting, we simulate 200 stages following the staged recruitment regime depicted in Figure 3 in Appendix A, and we repeat the full 200-stage simulation 50 times. Appendix B.1 provides details on the setup and a link to our implementation.
Figure 1: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings. DML-TS-NNR performs competitively in the first setting (the simplest), and it substantially outperforms the next-best method (IntelPooling) in the others. Pairwise comparisons in Table 2 in Appendix B.3) reveal that DML-TS-NNR outperformed all competitors shown above in 49 of 50 repetitions in the Nonlinear setting (panel (c)).
Figure 1 displays the average cumulative regret for each method. DML-TS-NNR and IntelPooling have similar performance in the Homogeneous Users setting, but DML-TS-NNR excels in the Heterogeneous Users and Nonlinear settings, with significantly lower cumulative regret by stage 200. This is expected as DML-TS-NNR can utilize network information and model nonlinear baselines in these settings. However, in the Homogeneous Users setting, where no network information is available and a linear baseline is used, DML-TS-NNR does not outperform IntelPooling.
Appendices B.2, B.3, and B.4 offer more results, including other method comparisons, a rectangular data array simulation, and pairwise statistical comparisons. The latter show DML-TS-NNR outperformed all competitors in at least 49 of 50 repetitions in the Nonlinear setting, indicating even higher statistical confidence than Figure 1 suggests.
@@ 6.2 Valentine Study Analysis Results
In this section, we compare DML-TS-NNR to the above algorithms via off-policy evaluation using data from the Valentine Study (Jeganathan et al., 2022), a prospective, randomized-controlled, remotely-administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients. In the analyzed dataset, participants were randomized to receive or not receive contextually tailored notifications promoting low-level physical activity and exercise throughout the day. The left of Figure 2 shows the estimated improvement in the average reward over the original constant randomization, averaged over stages (K = 120) and participants (N=108). We see that DML-TS-NNR achieved the highest average reward; its average performance (across bootstrap replications) is higher than the 75th percentile of all other methods.
To further analyze the improvement of DML-TS-NNR relative to the other methods, we aim to test whether the proposed algorithm consistently and significantly improves cumulative rewards across all assessed datasets. Here we conducted pairwise paired t-tests with one-sided alternative hypotheses. The null hypothesis ($H_{0}$) for these tests is that the two algorithms being compared achieve the same average reward. The alternative hypothesis ($H_{1}$) is that the algorithm listed in the column achieves higher average rewards than the algorithm listed in the row. The right of Figure 2 displays the p-values obtained from these pairwise t-tests. The last column’s dark shade indicates that the proposed DML-TS-NNR algorithm achieves significantly higher rewards than the other five competing algorithms ($p\leq 0.01$ for all pairwise comparisons).. This implies that after implementing our proposed algorithm, the step counts increased by $3.5\%$ more than what the constant randomization policy achieved. This translates to an increase of 140 steps per day in step count, given a baseline of 1000 steps per hour. See Appendix C for additional implementation details.
We performed an additional off-policy comparison using data from the Intern Health Study (IHS) (NeCamp et al., 2020), as shown in Figure 10. The results further demonstrate the competitive performance of our DML-TS-NNR algorithm, with the AC algorithm also showing comparable performance in this dataset. Further details on the analysis can be found in Appendix D.
@@ 7 Discussion
This paper introduces DML-TS-NNR, a new contextual bandit algorithm for mHealth studies, using DML and network cohesion penalties to model complex rewards and pool information across individuals and time. This enhances statistical precision, enabling rapid learning of accurate, context-specific mHealth treatment policies and improved rewards.
While DML-TS-NNR surpasses current methods, it has four potential improvement areas: it only considers immediate rewards, ignoring treatment fatigue; its computation of log-determinants and matrix inverses is expensive for large matrices; it assumes linear differential rewards and a known network with binary edges; and it requires specification of several hyperparameters. Future work will address these issues.
Appendix A Recruitment Regime Illustration
Figure 3: Illustration of the staged recruitment scheme. At each recruitment stage (each time point), a new participant is recruited and observed; at the same time, all participants who were recruited prior to the current stage are also observed again. Observations are not collected from participants who have yet to be recruited. For simplicity, we assume one participant is recruited at each stage.
Appendix B Additional Details for Simulation Study
@@
DML-TS-NNR $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
DML-TS-NNR-BLM $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$
DML-TS-SU $\times$ $\surd$ $\surd$ N/A $\surd$ $\surd$ $\surd$
@@
Table 1: This table compares the methods tested in the simulation study based on the design components listed in the columns, with $\surd$ indicating that the method includes the component listed in the column, $\times$ indicating that it does not, and N/A indicating that the component is not applicable. This table can be used to investigate the effect of certain model components on cumulative regret; for example, comparing DML-TS-NNR and DML-TS-NNR-BLM helps us understand the impact of modeling the nonlinear baseline—the only column that differs between these two methods.
B.1 Setup Details
The code for the simulation study is fully containerized and publicly available at https://github.com/anonymous-conf-submitter-896697/mixed-effects-bandit. The simulation study involves a generative model of the following form for user $i$ at time $t$:
$R_{it}=g(S_{it})+x(S_{it},A_{it})^{{\top}}\,\theta_{it}+\epsilon_{it},\quad\epsilon_{it}\sim\mathcal{N}(0,1)$
Here $S_{it}=(s_{1},s_{2})\in\mathbb{R}^{2}$ is a context vector, with both dimensions $\overset{iid}{\sim}U(-1,1)$. We set $x(s,a)=a\,(1,s_{1},s_{2})$. For simplicity, we set $g$ to a time-homogeneous function. The specific nature of the function varies across the following three settings mentioned in Section 6.1:
  • Homogeneous Users: Standard contextual bandit assumptions with a linear baseline and no user- or time-specific parameters. The linear baseline is $g(S_{it})=2-2s_{1}+3s_{2}$, and the causal parameter is $\theta_{it}=(1,0.5,-4)$ such that the optimal action varies across the state space.
@@ B.1 Setup Details
Figure 4: (left) The baseline reward function $g(S_{it})$ used in the simulation study. The proposed method allows this function to be a nonlinear function of the context vectors. The baseline was generated using a combination of recursive partitioning and by summing scaled, shifted, and rotated Gaussian densities. (right) The time-specific parameters used in the simulation study. These parameters cause the advantage function to vary over time. We set them such that the advantage function changes quickly at the beginning of the study then stabilizes.
We assume that the data are observed via a staged recruitment scheme, as illustrated in Figure 3 in Appendix A. For computational convenience, we update parameters and select actions in batches. If, for instance, we observe twenty users at a given stage, we update our estimates of the relevant causal parameters and select actions for all twenty users simultaneously. This strategy offers a slight computational advantage with limited implications in terms of statistical performance.
For simplicity, we assume that the nearest-neighbor network is known and set the relevant hyperparameters accordingly. We took care to set the hyperparameters such that DML-TS-NNR performs a similar amount of shrinkage compared to other methods, especially IntelPooling, which effectively uses a separate penalty matrix for users and time. To accomplish this, we used the same penalty matrices for DML-TS-NNR and IntelPooling, effectively generalizing the scalar $\gamma$ penalty discussed in the main paper. For the purposes of calculating the constant $B$ given in Algorithm 2, we used a separate value of $\gamma$ for both users and time and set them equal to the maximum eigenvalue of the corresponding penalty matrix. We use 5 neighbors within the DML methods and set the other hyperparameters as follows: $\lambda=1$, and $\delta=0.01$, and $v_{k}=1$ (i.e., $v_{k}$ is constant).
@@
Figure 5: (left) The baseline reward function $g(S_{it})$ used in the simulation study compared to (right) the estimated baseline reward from our neural network in the nonlinear setting.
@@
We include the Neural-Linear method primarily to demonstrate that correctly modeling the baseline is not sufficient to ensure good performance. In the mobile health context, algorithms should also be able to (1) efficiently pool data across users and time and (2) leverage network information. The Neural-Linear method satisfies neither of these criteria. Note that a neural network could be used to model the baseline rewards as part of our algorithm. Future work could consider allowing the differential rewards themselves to also be complex nonlinear functions, which could be accomplished by combining our method with Neural-Linear. We leave the details to future work.
B.2 Additional Method Comparisons
@@
Our results in the main paper indicate that DML-TS-NNR outperforms existing methods, especially in the complex longitudinal setting that we are targeting. These positive results beg the question: What aspects of DML-TS-NNR contribute most to its superior performance?
@@
In this section, we describe an additional simulation study designed to answer this question. In it, we compare DML-TS-NNR (as implemented in the main paper) to three additional comparison methods: DML-TS-NNR-BLM, DML-TS-SU, and NNR-Linear. Each comparison isolates particular aspects of our algorithm to understand its contribution to DML-TS-NNR’s performance.
@@
DML-TS-NNR-BLM is a version of DML-TS-NNR that employs Bagged Linear Models as the baseline supervised learning algorithm. Unlike the method implemented in the main paper (which uses bagged stochastic gradient trees), this baseline model is misspecified in the nonlinear setting; i.e., it cannot accurately model the nonlinear baseline reward function. Consequently, this comparison will allow us to gauge the impact of the supervised learning method used in our method. In particular, it should allow us to validate whether the algorithm is robust to baseline misspecification as the theory suggests.
@@
DML-TS-SU is a Single-User version of our algorithm. Rather than having a separate $\theta_{i}$ for each user, this method estimates a single advantage function that is shared across all users. It does, however, employ DML, time effects, and nearest-neighbor regularization (for the time effects). Consequently, this comparison will help us understand the impact of the user-specific parameters in our simulation.
@@
NNR-Linear is similar to the existing network-cohesion bandit algorithms described in the main paper. It includes a distinct $\theta_{i}$ for each user and regularizes them toward each other using a Laplacian penalty. It does not, however, use DML or time effects. As a result, this comparison isolates the impact of these two factors and should provide evidence that DML-TS-NNR is more broadly applicable than existing network-cohesion bandit algorithms.
Figure 6 displays the cumulative regret as a function of the stage for these methods across the three settings used in the main paper. We see that all methods perform similarly in panel (a), the setting with homogeneous users. However, in panel (b) we see that the regret achieved by DML-TS-SU begins to resemble a straight upward-sloping line because it cannot appropriately model the user-specific advantage functions.
In panel (c), we observe a much wider spread of performance, with DML-TS-NNR and DML-TS-NNR-BLM performing similarly and substantially outperforming the other methods. The fact that DML-TS-NNR and DML-TS-NNR-BLM perform similarly indicates that our algorithm is, in fact, robust to misspecification as expected. The comparison to NNR-Linear reveals that the DML and time effects—the new components of our method compared to existing network-cohesion bandit algorithms—do contribute meaningfully to DML-TS-NNR’s superior performance. We found in additional simulations studies not reported here that the gap between NNR-Linear and our method widens as the magnitude of the time effects increases.
@@
Figure 6: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings for the additional comparison methods. The fact that DML-TS-NNR outperforms DML-TS-SU demonstrates the importance of the user-specific parameters. The fact that DML-TS-NNR outperforms NNR-Linear shows that previous network cohesion approaches do not adequately address the (1) nonlinear nature of the baseline reward and (2) presence of time effects. The comparable performance of DML-TS-NNR-BLM to DML-TS-NNR highlights how our algorithm is robust to misspecification of the baseline reward model.
B.3 Pairwise Comparisons

### v4

@@ 6 Experiments
This section presents results from applying RoME in a simulated mHealth study with staggered recruitment and an off-policy comparison study for cardiac rehabilitation mHealth interventions. The simulations were implemented using Python and the results were generated using individual compute nodes with two 3.0 GHz Intel Xeon Gold 6154 processors and 180 GB of RAM. Case studies were implemented using R 4.2.2 and results were generated on a cluster composed of individual compute nodes with 2.10 GHz Intel Xeon Gold 6230 processors and 192 GB of RAM.
@@ 6.1 Competitor Comparison Simulation
In this section, we compare RoME to four competing methods in simulation. We implemented RoME using Option 2 with a bagged ensemble of stochastic gradient trees (Gouk et al., 2019; Mastelini et al., 2021) trained online via the River library (Montiel et al., 2021). We programmed the linear algebra operations using SuiteSparse (Davis & Hu, 2011) to take advantage of the sparse nature of the Laplacian matrices and $\phi_{i,t}$ from Algorithm 2.
@@
We compare these methods under three settings: Homogeneous Users, Heterogeneous Users, and Nonlinear. The first two involve a linear baseline model and time-homogeneous parameters, with the second setting having distinct user parameters. The third setting, designed to mirror the challenges of mHealth studies, includes a nonlinear baseline and both user- and time-specific parameters. For each setting, we simulate 200 stages following the staged recruitment regime depicted in Figure 1, and we repeat the full 200-stage simulation 50 times. Appendix A.1 provides details on the setup and a link to our implementation.
Figure 2: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings. RoME performs competitively in the first setting (the simplest), and it substantially outperforms the next-best method (IntelPooling) in the others.
Figure 2 displays the average cumulative regret for each method. RoME and IntelPooling have similar performance in the Homogeneous Users setting, but RoME excels in the Heterogeneous Users and Nonlinear settings, with significantly lower cumulative regret by stage 200. This is expected as RoME can utilize network information and model nonlinear baselines in these settings. However, in the Homogeneous Users setting, where no network information is available and a linear baseline is used, RoME does not outperform IntelPooling.
Appendices A.2A.5 offer more results, including additional method comparisons, a rectangular data array simulation, hyperparameter sensitivity analyses, and pairwise statistical comparisons. The latter show that RoME outperformed each competitor in at least 48 of 50 repetitions in the Nonlinear setting, indicating even higher statistical confidence than Figure 2 suggests. RoME substantially outperforms the competing algorithms in the additional comparisons and is several orders of magnitude faster than would be required in a standard mHealth study, producing over 20,000 decisions in as little as 95 seconds (see Table 2 in Appendix A.2).
@@ 6.2 Valentine Study Analysis Results
In this section, we compare RoME to the above algorithms via off-policy evaluation using data from the Valentine Study (Jeganathan et al., 2022), a prospective, randomized-controlled, remotely administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients. In the analyzed dataset, participants were randomized to receive or not receive contextually tailored notifications promoting low-level physical activity and exercise throughout the day. The left panel of Figure 3 shows the estimated improvement in the average reward over the original constant randomization, averaged over stages (K = 120) and participants (N=108). We see that RoME achieved the highest average reward; its average performance (across bootstrap replications) is higher than the 75th percentile of all other methods.
To further analyze the improvement of RoME relative to the other methods, we test whether the proposed algorithm significantly improves cumulative rewards using paired t-tests with one-sided alternative hypotheses. The null hypothesis ($H_{0}$) for these tests is that the two algorithms being compared achieve the same average reward. The alternative hypothesis ($H_{1}$) is that the algorithm listed in the column achieves higher average rewards than the algorithm listed in the row. The right of Figure 3 displays the p-values obtained from these pairwise t-tests. The dark shade of the last column indicates that the proposed RoME algorithm achieves significantly higher rewards than the other five competing algorithms ($p\leq 0.01$ for all pairwise comparisons). The results imply that RoME would have achieved a $3.5\%$ increase in step count relative to the constant randomization policy that was actually implemented in this study. This effect translates to an increase of 140 steps per day, given a baseline of 1,000 steps per hour and four one-hour measurement windows. See Appendix B for additional implementation details.
We performed an additional off-policy comparison using data from the Intern Health Study (IHS) (NeCamp et al., 2020), as shown in Figure 11 in Appendix C.2. The results further demonstrate the competitive performance of the RoME algorithm, with the AC algorithm also showing comparable performance in this dataset. Further details on the analysis can be found in Appendix C.
@@ 7 Discussion
This paper introduces RoME, a robust mixed-effects contextual bandit algorithm for mHealth studies. RoME adapts DML and network cohesion penalties to dynamic settings, enabling researchers to efficiently pool information across users and over time. In addition to the methodological contribution, we also prove a high-probability regret bound in an asymptotic regime that involves a growing pool of users and time points. Our implementation of RoME is several orders of magnitude faster than would be required in practice and could easily be adapted for use in future mHealth studies.
We see several promising directions for improving RoME in future work. One such direction might consider improvements in the algorithm and theoretical analysis that enable us to replace Assumptions 8 and 9 with weaker and more primitive assumptions. In particular, these assumptions may not be plausible for $q>1$ because most arm selections will be restricted to the same two arms (for large $K$). Other interesting directions include data-adaptive strategies for hyperparameter selection and addressing long-term impacts of mHealth treatments, such as accumulating treatment fatigue. Future work could also consider computational improvements that would enable large-scale deployment of RoME, which would be especially relevant in nonclinical settings.
Acknowledgements
The authors gratefully acknowledge the contributions of the researchers, administrators, and participants involved in the Valentine (https://clinicaltrials.gov/study/NCT04587882) and Intern Health (https://clinicaltrials.gov/study/NCT03972293) studies. This work was supported in part by grant P50 DA054039 provided through the NIH and NIDDK and by grant R01 GM152549 through the NIH and NIGMS. Dr. Golbus receives funding from the NIH (L30HL143700, 1K23HL168220). Initially while working on this paper, Dr. Moreno was supported by Luminous Computing, Inc.
Appendix A Additional Details for Simulation Study
@@
RoME $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
RoME-BLM $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$
RoME-SU $\times$ $\surd$ $\surd$ N/A $\surd$ $\surd$ $\surd$
@@
Table 1: This table compares the methods tested in the simulation study based on the design components listed in the columns, with $\surd$ indicating that the method includes the component listed in the column, $\times$ indicating that it does not, and N/A indicating that the component is not applicable. This table can be used to investigate the effect of certain model components on cumulative regret; for example, comparing RoME and RoME-BLM helps us understand the impact of modeling the nonlinear baseline—the only column that differs between these two methods.
A.1 Setup Details
The code for the simulation study is fully containerized and publicly available at https://github.com/eastonhuch/RoME. The simulation study involves a generative model of the following form for user $i$ at time $t$:
$R_{it}=g(S_{i,t})+x(S_{i,t},A_{i,t})^{{\top}}\,\theta_{it}+\epsilon_{it},\quad\epsilon_{it}\sim\mathcal{N}(0,1)$
Here $S_{i,t}=(s_{1},s_{2})\in\mathbb{R}^{2}$ is a context vector, with both dimensions $\overset{iid}{\sim}U(-1,1)$. We set $x(s,a)=a\,(1,s_{1},s_{2})$. For simplicity, we set $g$ to a time-homogeneous function. The specific nature of the function varies across the following three settings mentioned in Section 6.1:
  • Homogeneous Users: Standard contextual bandit assumptions with a linear baseline and no user- or time-specific parameters. The linear baseline is $g(S_{i,t})=2-2s_{1}+3s_{2}$, and the causal parameter is $\theta_{it}=(1,0.5,-4)$ such that the optimal action varies across the context space.
@@ A.1 Setup Details
Figure 4: (left) The baseline reward function $g_{t}(S_{i,t})$ (constant across time in this case) used in the simulation study. The proposed method allows this function to be a nonlinear function of the context vectors. The baseline was generated using a combination of recursive partitioning and by summing scaled, shifted, and rotated Gaussian densities. (right) The time-specific parameters used in the simulation study. These parameters cause the advantage function to vary over time. We set them such that the advantage function changes quickly at the beginning of the study then stabilizes.
We assume that the data are observed via a staged recruitment scheme, as illustrated in Figure 1. For computational convenience, we update parameters and select actions in batches. If, for instance, we observe twenty users at a given stage, we update our estimates of the relevant causal parameters and select actions for all twenty users simultaneously. This strategy offers a slight computational advantage with limited implications in terms of statistical performance.
For simplicity, we assume that the nearest-neighbor network is known and set the relevant hyperparameters accordingly. We took care to set the hyperparameters such that RoME performs a similar amount of shrinkage compared to other methods, especially IntelPooling, which effectively uses a separate penalty matrix for users and time. To accomplish this, we used the same penalty matrices for RoME and IntelPooling, effectively generalizing the scalar $\gamma$ penalty discussed in the main paper. We use 5 neighbors within the DML methods and set the other hyperparameters as follows: $\lambda=1$, $\delta=0.01$, $v=1$, and $\zeta=10$.
@@
Figure 5: (left) The baseline reward function $g_{t}(S_{i,t})$ used in the simulation study compared to (right) the estimated baseline reward from our neural network in the nonlinear setting.
@@
We include the Neural-Linear method primarily to demonstrate that correctly modeling the baseline is not sufficient to ensure good performance. In the mHealth, algorithms should also be able to (a) efficiently pool data across users and time and (b) leverage network information. The Neural-Linear method satisfies neither of these criteria. Note that a neural network could be used to model the baseline rewards as part of our algorithm. Future work could consider allowing the differential rewards themselves to also be complex nonlinear functions, which could be accomplished by combining our method with Neural-Linear. We leave the details to future work.
A.2 Additional Method Comparisons
@@
Our results in the main paper indicate that RoME outperforms existing methods, especially in the complex longitudinal setting that we are targeting. These positive results beg the question: What aspects of RoME contribute most to its superior performance?
@@
In this section, we describe an additional simulation study designed to answer this question. In it, we compare RoME (as implemented in the main paper) to three additional comparison methods: RoME-BLM, RoME-SU, and NNR-Linear. Each comparison isolates particular aspects of our algorithm to understand its contribution to RoME’s performance.
@@
RoME-BLM is a version of RoME that employs Bagged Linear Models as the baseline supervised learning algorithm. Unlike the method implemented in the main paper (which uses bagged stochastic gradient trees), this baseline model is misspecified in the nonlinear setting; i.e., it cannot accurately model the nonlinear baseline reward function. Consequently, this comparison will allow us to gauge the impact of the supervised learning method used in our method. In particular, it should allow us to validate whether the algorithm is robust to baseline misspecification as the theory suggests.
@@
RoME-SU is a Single-User version of our algorithm. Rather than having a separate $\theta_{i}$ for each user, this method estimates a single advantage function that is shared across all users. It does, however, employ DML, time effects, and nearest-neighbor regularization (for the time effects). Consequently, this comparison will help us understand the impact of the user-specific parameters in our simulation.
@@
NNR-Linear is similar to the existing network-cohesion bandit algorithms described in the main paper. It includes a distinct $\theta_{i}$ for each user and regularizes them toward each other using a Laplacian penalty. It does not, however, use DML or time effects. As a result, this comparison isolates the impact of these two factors and should provide evidence that RoME is more broadly applicable than existing network-cohesion bandit algorithms.
@@
Figure 6: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings for the additional comparison methods. The fact that RoME outperforms RoME-SU demonstrates the importance of the user-specific parameters. The fact that RoME outperforms NNR-Linear shows that previous network cohesion approaches do not adequately address the nonlinear nature of the baseline reward and presence of time effects. The comparable performance of RoME-BLM to RoME highlights how our algorithm is robust to misspecification of the baseline reward model.
Figure 6 displays the cumulative regret as a function of the stage for these methods across the three settings used in the main paper. We see that all methods perform similarly in panel (a), the setting with homogeneous users. However, in panel (b) we see that the regret achieved by RoME-SU begins to resemble a straight upward-sloping line because it cannot appropriately model the user-specific advantage functions.
In panel (c), we observe a much wider spread of performance, with RoME and RoME-BLM performing similarly and substantially outperforming the other methods. The fact that RoME and RoME-BLM perform similarly indicates that our algorithm is, in fact, robust to misspecification as expected. The comparison to NNR-Linear reveals that the DML and time effects—the new components of our method compared to existing network-cohesion bandit algorithms—do contribute meaningfully to RoME’s superior performance. We found in additional simulations studies not reported here that the gap between NNR-Linear and our method widens as the magnitude of the time effects increases.
Homogeneous Users Heterogeneous Users Nonlinear
RoME $394.5$ $406.4$ $420.1$
RoME-BLM $94.6$ $96.8$ $109.9$
RoME-SU $383.5$ $390.5$ $396.5$
NNR-Linear $29.7$ $31.3$ $31.5$
IntelPooling $46.6$ $42.1$ $45.0$
Neural-Linear $0.7$ $0.8$ $0.9$
Standard $0.2$ $0.2$ $0.3$
AC $0.1$ $0.1$ $0.3$
Table 2: Average computation time across all methods and settings in the main simulation.
Table 2 displays the average computation time for each method across all three settings. Each run of the simulation produces $200\times 201/2=20,100$ decisions. RoME and RoME-SU require the longest computation time at about seven minutes. RoME-BLM requires only 95–110 seconds, indicating that the computational bottleneck for RoME is the ML model fitting. The required computation time for IntelPooling is about half that of RoME-BLM at 42–47 seconds. NNR-Linear requires 30–32 seconds, and the remaining methods run in less than one second. Because most clinical mHealth studies require fewer than 1,000 decisions per day, the speed of RoME is orders of magnitude faster than is required in practice. However, large-scale commercial deployment of RoME may require computational adjustments and approximations.
A.3 Pairwise Comparisons